Find the time required to liberate 1.0 litre of hydrogen at STP in an electrolytic cell by a current of 5.0 A.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. H base 2/22.4L → 2g 1L→2/22.4 M = Zit 2/22.4 = 1/96500 *5*T ⇒ T = 2/22.4*96500/5 = 1732.21 sec = 28.7 min = 29 min.